A load of 10 kg is suspended by a metal wire 3 m long and having a cross-sectional area 4 mm2. Find (a) the stress (b) the strain and (c) the elongation. Young modulus of the metal is 2.0 × 1011 N m-2.

Grade:upto college level

2 Answers

Jitender Pal
9 years ago
Sol. F = mg Stress = F/A Strain = ∆L/L Y = FL/A∆L ⇒ ∆L/L = F/YA
Ravi
5 years ago
Load 12kg wire 3m area 5mm2.ccacalcalccalcucalculcalculate sststrstrestreastress sststrstrastranstrastraistrainstrain